OpenStudy (loser66):

Suppose \(|f(x)|\leq M\) for all \(x\in[a,b]\). For \(x\in[a,b]\) set \(F(x)=\int_a^x f(t)dt\). Show that F(x) is continuous on [a,b] Please, help

OpenStudy (loser66):

@GIL.ojei

OpenStudy (anonymous):

for a function to be continues, what must he posses?

OpenStudy (loser66):

\(lim_{x\rightarrow x_0} f(x) = f(x_0)\)

OpenStudy (anonymous):

very good. it must posses a limit. or its limit must exist

OpenStudy (loser66):

for all x_0 on [a,b]

OpenStudy (anonymous):

yap

OpenStudy (loser66):

then?

OpenStudy (anonymous):

\[F(x)=\int\limits_a^x f(t)dt \] that is the function. integrate it with respect to t and subtitute a and x and conclude

OpenStudy (anonymous):

using \[ |f(x)|≤M for all x∈[a,b] \] to solve

OpenStudy (loser66):

Got you

OpenStudy (loser66):

I think of Riemann integrable of F(x) to show that F(x) integrable, hence it is continuous. But it seems I complicate the problem

OpenStudy (anonymous):

hmmm

OpenStudy (loser66):

ok, let finish it. :) to make sure I don't overestimate my ability. hehehe

OpenStudy (anonymous):

can you intigrate\[F(x)=\int\limits_a^x f(t)dt \]

OpenStudy (loser66):

why not? it just f(x) - f(a) right?

OpenStudy (loser66):

and f(x) - f(a) < f(x) |f(x) -f(a) |<|f(x) < M

OpenStudy (loser66):

I will show my stuff here. If you can, check when you are back, please. Thanks for the help though

OpenStudy (helder_edwin):

u have to use the definition of continuity: \[\large (\forall\varepsilon>0)(\exists\delta>0)(\forall x)(|x-x_0|<\delta\Rightarrow |F(x)-F(x_0)|<\varepsilon ) \]

OpenStudy (helder_edwin):

Let \(c\in[a.b]\). If \(h>0\) then \[\large F(c+h)-F(c)=\int_a^{c+h}f-\int_a^cf=\int_c^{c+h}f. \]

OpenStudy (helder_edwin):

Since \(-M\leq f(x)\leq M\), then \[ -M\cdot h\leq\int_c^{c+h}f\leq M\cdot h \] in other words \[ -M\cdot h\leq F(c+h)-F(c)\leq M\cdot h.\tag{1} \] If \(h<0\) u can get a similar inequality.

OpenStudy (helder_edwin):

Notice that \[ F(c+h)-F(c)=\int_c^{c+h}f=-\int_{c+h}^cf. \] Then, over the interval \([c+h,c]\) of length \(-h\) we have \[ M\cdot h\leq\int_{c+h}^cf\leq-M\cdot h \] which when multiplied by \(-1\) becomes \[ M\cdot h\leq F(c+h)-F(c)\leq-M\cdot h\tag{2} \]

OpenStudy (helder_edwin):

(1) and (2) can be combined into \[ |F(c+h)-F(c)|\leq M\cdot|h|. \] Therefore, if \(\varepsilon>0\), we have that \[ |F(c+h)-F(c)|<\varepsilon, \] whenever \(|h|<\varepsilon/M\). This proves that \[ \lim_{h\to0}F(c+h)=F(c) \] which means that \(F\) is continous at \(c\).

OpenStudy (anonymous):

now this is very corrct. thanks @helder_edwin

OpenStudy (loser66):

Thanks for the help. Much appreciate. @GIL.ojei @helder-edwin